OpenStudy (anonymous):

how to find the zeros of (1/3)x^3(x-4)^2

OpenStudy (cwrw238):

by inspection x = 0 is one root as 1/3 x^3 = 0

OpenStudy (cwrw238):

also (x-4)^2 = 0 so what does that make x?

OpenStudy (cwrw238):

the concept is if 2 functions multiplied by each other = zero then either can be zero (1/3)x^3 = 0 or (x-4)^2 = 0